OpenStudy (anonymous):

Evaluate the following using exact radian measure Sin-1(-3squared/2)

jimthompson5910 (jim_thompson5910):

Did you mean \[\Large \sin^{-1}\left(-\frac{\sqrt{3}}{2}\right)\] ???

OpenStudy (anonymous):

Yes!

jimthompson5910 (jim_thompson5910):

locate the point in Q4 that has a y coordinate of \[\Large -\frac{\sqrt{3}}{2}\]

jimthompson5910 (jim_thompson5910):

the corresponding angle will be the answer

jimthompson5910 (jim_thompson5910):

well not quite, but anyways, tell me what angle you get

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

I seriously don't know what I'm doing

jimthompson5910 (jim_thompson5910):

are you able to find the point that has a y coordinate of -sqrt(3)/2 ?

jimthompson5910 (jim_thompson5910):

use that unit circle I gave you